OpenStudy (anonymous):

I need help! : Solve each of the following. Please give both the general solution and the specific solutions in the given interval. 2 sin (2 theta) + sqrt3 = 0. intervals (0,2pi)

jimthompson5910 (jim_thompson5910):

First step?

OpenStudy (anonymous):

Hmm...... im not totally sure

jimthompson5910 (jim_thompson5910):

We want to isolate theta. So we first move that square root of 3 term over by subtracting it from both sides. \[\Large 2\sin\left(2\theta\right) + \sqrt{3} = 0\] \[\Large 2\sin\left(2\theta\right) = -\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

What's next?

OpenStudy (anonymous):

Ok then would we divide by 2...?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

\[\Large 2\sin\left(2\theta\right) + \sqrt{3} = 0\] \[\Large 2\sin\left(2\theta\right) = -\sqrt{3}\] \[\Large \sin\left(2\theta\right) = -\frac{\sqrt{3}}{2}\]

jimthompson5910 (jim_thompson5910):

Next?

OpenStudy (anonymous):

we...divide by sin? or 2?

OpenStudy (anonymous):

im not really sure

jimthompson5910 (jim_thompson5910):

that's a good guess, but that's not it

jimthompson5910 (jim_thompson5910):

sine is a function

jimthompson5910 (jim_thompson5910):

so we need to apply the inverse function to get rid of it

jimthompson5910 (jim_thompson5910):

the inverse of sine is either called "inverse sine" or "arcsine"

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

so take the arcsine of both sides... \[\Large 2\sin\left(2\theta\right) + \sqrt{3} = 0\] \[\Large 2\sin\left(2\theta\right) = -\sqrt{3}\] \[\Large \sin\left(2\theta\right) = -\frac{\sqrt{3}}{2}\] \[\Large 2\theta = \arcsin\left(-\frac{\sqrt{3}}{2}\right)\]

OpenStudy (anonymous):

Ok cool, thats what i was thinking

OpenStudy (anonymous):

then what would be the nxt step?

jimthompson5910 (jim_thompson5910):

Now use this unit circle http://www.math.ucsd.edu/~jarmel/math4c/Unit_Circle_Angles.png to evaluate the arcsine of \(\Large -\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

is it 4pi/3?

OpenStudy (anonymous):

and 5pi/3

OpenStudy (anonymous):

can you explain how to find that?

jimthompson5910 (jim_thompson5910):

You found the correct angles, you found the angles x such that \[\Large \sin(x) = -\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

Ok what do I do then, like how to formulate that info into the equation?

jimthompson5910 (jim_thompson5910):

You can A) Use the unit circle B) Memorize the angles C) Use a calculator to find the arcsine of a value

jimthompson5910 (jim_thompson5910):

I like option A because it allows for exact answers and you don't have to memorize anything.

OpenStudy (anonymous):

ok what would be the nexts tep

jimthompson5910 (jim_thompson5910):

So you found that the arcsine of \(\Large -\frac{\sqrt{3}}{2}\) is 4pi/3 or 5pi/3

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

This means we get... \[\Large 2\sin\left(2\theta\right) + \sqrt{3} = 0\] \[\Large 2\sin\left(2\theta\right) = -\sqrt{3}\] \[\Large \sin\left(2\theta\right) = -\frac{\sqrt{3}}{2}\] \[\Large 2\theta = \arcsin\left(-\frac{\sqrt{3}}{2}\right)\] \[\Large 2\theta = \frac{4\pi}{3} \ \text{or} \ 2\theta = \frac{5\pi}{3}\]

OpenStudy (anonymous):

Ok makes sense

jimthompson5910 (jim_thompson5910):

Now because any coterminal angle to 4pi/3 or 5pi/3 works, this means that we can add multiples of 2pi to each angle

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

\[\Large 2\sin\left(2\theta\right) + \sqrt{3} = 0\] \[\Large 2\sin\left(2\theta\right) = -\sqrt{3}\] \[\Large \sin\left(2\theta\right) = -\frac{\sqrt{3}}{2}\] \[\Large 2\theta = \arcsin\left(-\frac{\sqrt{3}}{2}\right)\] \[\Large 2\theta = \frac{4\pi}{3} \ \text{or} \ 2\theta = \frac{5\pi}{3}\] \[\Large 2\theta = \frac{4\pi}{3}+2\pi k \ \text{or} \ 2\theta = \frac{5\pi}{3}+2\pi k\]

jimthompson5910 (jim_thompson5910):

where k is any integer

jimthompson5910 (jim_thompson5910):

Last step?

OpenStudy (anonymous):

divid 2?

OpenStudy (anonymous):

so theta is isolated?

jimthompson5910 (jim_thompson5910):

yes, divide everything by 2 \[\Large 2\sin\left(2\theta\right) + \sqrt{3} = 0\] \[\Large 2\sin\left(2\theta\right) = -\sqrt{3}\] \[\Large \sin\left(2\theta\right) = -\frac{\sqrt{3}}{2}\] \[\Large 2\theta = \arcsin\left(-\frac{\sqrt{3}}{2}\right)\] \[\Large 2\theta = \frac{4\pi}{3} \ \text{or} \ 2\theta = \frac{5\pi}{3}\] \[\Large 2\theta = \frac{4\pi}{3}+2\pi k \ \text{or} \ 2\theta = \frac{5\pi}{3}+2\pi k\] \[\Large \theta = \frac{4\pi}{6}+\pi k \ \text{or} \ \theta = \frac{5\pi}{6}+\pi k\] \[\Large \theta = \frac{2\pi}{3}+\pi k \ \text{or} \ \theta = \frac{5\pi}{6}+\pi k\]

OpenStudy (anonymous):

ok is that my final?

jimthompson5910 (jim_thompson5910):

yes, those are your general solutions

OpenStudy (anonymous):

wait if you divide by 2 wouldnt the 5pi/6 change?

OpenStudy (anonymous):

and the 2pi/3??

jimthompson5910 (jim_thompson5910):

I did that. I started with 4pi/3 and divided by 2 to get 4pi/6. Then reduced to get 2pi/3

jimthompson5910 (jim_thompson5910):

Then I divided 5pi/3 by 2 to get 5pi/6

OpenStudy (anonymous):

Oh ok:) thx, you are really the best. do u mind if i ask 2 more? i can tell you what i wpuld do next so youre not doing everything? sound good ?

jimthompson5910 (jim_thompson5910):

wait a sec

jimthompson5910 (jim_thompson5910):

those were the general solutions

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

they now want the "specific solutions in the given interval"

jimthompson5910 (jim_thompson5910):

and the given interval is (0,2pi)

OpenStudy (anonymous):

Ok, im not 100% sure how to find that

jimthompson5910 (jim_thompson5910):

So how do we find these solutions?

OpenStudy (anonymous):

the specific

OpenStudy (anonymous):

hmmm.......

jimthompson5910 (jim_thompson5910):

Here's the best way to do find the solutions in the given interval Start with \[\Large \theta = \frac{2\pi}{3}+\pi k \] and plug in k = 0 What do you get?

OpenStudy (anonymous):

I get theta=2pi/3 + 0

jimthompson5910 (jim_thompson5910):

good, is that in the interval (0,2pi) ???

OpenStudy (anonymous):

yes?

jimthompson5910 (jim_thompson5910):

you can check 2pi = 6.283185 2pi/3 = 2.094395 So 2.094395 is definitely between 0 and 6.283185

jimthompson5910 (jim_thompson5910):

Therefore, 2pi/3 is between 0 and 2pi

OpenStudy (anonymous):

ok!

jimthompson5910 (jim_thompson5910):

now plug in k = 1 and tell me what you get

OpenStudy (anonymous):

I get theta=2pi/3 + pi

jimthompson5910 (jim_thompson5910):

Ok, add those terms

OpenStudy (anonymous):

I got theta= 5pi/3

jimthompson5910 (jim_thompson5910):

good, now check to see if its in range

OpenStudy (anonymous):

it is i believe

OpenStudy (anonymous):

sry i dont have a calc near me right now

jimthompson5910 (jim_thompson5910):

5pi/3 = 5.23598775598299 which is less than 6.283185, so yes it is

jimthompson5910 (jim_thompson5910):

Now plug in k = 2 and simplify

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i get theta=8pi/3

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

8pi/3 = 8.37758040957279, is that in range?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

o r is it?

jimthompson5910 (jim_thompson5910):

you are correct, it's not in range

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so then what would my specific solution answers be?

jimthompson5910 (jim_thompson5910):

So far, we found that k = 0 and k = 1 work (but k = 2 does not), correct?

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

these k values produce theta values of 2pi/3 and 5pi/3

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

so those two solutions are specific solutions (so far..) Now we must check \[\Large \theta = \frac{5\pi}{3}+\pi k \]

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

If you plug in k = 0, you get 5pi/3 if you plug in k = 1, you get 8pi/3 (I'm sure you see the pattern) So the solutions overlap...which means that there are only 2 solutions in this interval

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

and those two solutions are 2pi/3 and 5pi/3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 2pi/3 and5pi/3 are sp[ecific soliutions

OpenStudy (anonymous):

specific solutions

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Ok! do u not mind helping with just two more?

jimthompson5910 (jim_thompson5910):

they are specific solutions to \(\Large 2\sin\left(2\theta\right) + \sqrt{3} = 0\) in the interval \(\Large \left(0,2\pi\right)\)

jimthompson5910 (jim_thompson5910):

sure, go for it

OpenStudy (anonymous):

Ok here is the next one: same directions as the first. the problem is: 2 sin^2 x = 2 + cos x --> interval (0,pi)

jimthompson5910 (jim_thompson5910):

ok, do you know any trig identities?